# Improved Fine-Tuning of Large Multimodal Models for Hateful Meme Detection

### v1

@@ 4.1 Comparing LMM-RGCL to Baseline Systems under Supervised Settings
13 LMM-RGCL + RKC 77.1 69.3 88.8 81.7 81.4 75.6 64.5 66.4 63.8 55.6 69.3 69.3
Table 2: Comparing out-of-domain meme classification performance under low-resource settings. For systems fine-tuned under cross-dataset settings, models are fine-tuned on HarMeme and evaluated on the HatefulMemes dataset. For the remaining evaluation datasets, models are fine-tuned on the HatefulMemes dataset. For LMMs, we compare the SFT models using zero-shot and in-context learning with the LMM-RGCL fine-tuned models using RKC. Few-shot examples (4-shot) and RKC examples are drawn from the training split of each evaluation dataset. #2 and #3 are taken from the original paper. Best performance is highlighted in bold; second-best is underlined.
@@
As shown in Table 1, RGCL (#4) achieves the highest performance among CLIP-based classifiers, surpassing standard fine-tuned CLIP (#1) by approximately 10% across multiple datasets. On 5 out of 6 datasets, RGCL performs better than, or on par with, all three SFT LMMs (#7, #11, #15).
@@
We compare the zero-shot (#5, #9, #13) and few-shot (#6, #10, #14) performance of the pre-trained LMMs. Our findings indicate that, unlike in some other tasks, in-context learning does not benefit meme classification, which is consistent with previous results Hee et al. (2024); Huang et al. (2024). HarMeme is the only dataset where few-shot systems consistently outperform zero-shot systems. On Harm-P and MultiOFF, although the accuracies of zero-shot and few-shot remain comparable, the few-shot experiments yield a significant gain in F1 score. This improvement is due to a more balanced precision and recall after providing demonstration examples to the system.
Observation 3: LMM-RGCL outperforms all strong baseline systems across six datasets
For the six datasets and the three LMMs, fine-tuning with LMM-RGCL significantly improves performance over SFT (Table 1: #7, #8; #11, #12; #15, #16). Statistical significance tests between LMM-RGCL and SFT further confirm this, with all p-values below 0.05. Notably, as indicated in #16, Qwen2VL-7B fine-tuned with LMM-RGCL outperforms VPD-PaLI-X-55B on HatefulMemes. Moreover, LMM-RGCL improves upon RGCL with gains of over 4% in AUC and 3% in accuracy on HatefulMemes. These gains show LMM-RGCL’s effectiveness in improving LMMs for meme classification over SFT.
4.2 Comparing LMM-RGCL with Baseline Systems under Low-Resource Settings
Online hate speech is constantly evolving, posing a challenge to systems as the distribution of memes encountered in the wild departs from that of the training data. To simulate real-world deployment constraints, we evaluate systems on out-of-domain examples under low-resource settings where gradient updates are prohibited and only demonstration examples are available Huang et al. (2024); Hee et al. (2024); Cao et al. (2024).
@@ 4.2 Comparing LMM-RGCL with Baseline Systems under Low-Resource Settings
We compare LMM-RGCL fine-tuned LMM with the RKC against the following systems: SFT LMMs with zero-shot and few-shot prompting using LMH; GPT-4o OpenAI (2024a); specialized low-resource systems (LOREHM Huang et al. (2024), Mod-hate Cao et al. (2024)). For GPT-4o, the token likelihood is not accessible to compute the AUC score. For Mod-Hate and LOREHM, we report the results from the original papers.
@@
Observation 2: In-context learning remains ineffective for SFT LMMs
As shown in Table 2 #6, #9 and #12, the few-shot approach remains similarly ineffective after LMMs are fine-tuned on different domains of hateful meme datasets, offering no significant gains over the SFT zero-shot models (Table 2 #5, #8, #11). In Section 4.4, we further compare the number of shots for in-context learning and find that adding more shots does not improve performance.
Observation 3: LMM-RGCL fine-tuned LMMs with RKC inference mode outperforms baseline methods
LMM-RGCL fine-tuned LMMs using RKC outperform the baseline SFT LMMs in both zero-shot and few-shot settings under the same cross-dataset fine-tuning settings. Notably, LMM-RGCL trained Qwen2VL-7B with RKC improves over the baseline SFT few-shot model by 21.6% in AUC and 19.3% in accuracy on HarMeme (Table 2 #11-13). The ablation study in Section 4.4, which varies the number of top k for RKC, further demonstrates that RKC uses demonstration examples more effectively than the few-shot in-context learning framework. Additionally, we compare RKC results for pre-trained and SFT LMMs in Appendix D.
Observation 4: LMM-RGCL trained LMMs with RKC inference outperform other low resource methods
Compared to GPT-4o, our LMM-RGCL fine-tuned Qwen2VL-7B with RKC achieves 14.5% higher accuracy on HarMeme and 11.3% higher accuracy on PrideMM. When compared to other low-resource methods using similar open-source LMMs, our LMM-RGCL fine-tuned LLaVA-1.5-7B with RKC matches the performance of LOREHM on the HatefulMemes dataset. Notably, LOREHM uses a much larger LLaVA-1.6-34B within an agent-based framework. Furthermore, our method outperforms LOREHM by 8.2% in accuracy on HarMeme, highlighting our methods’ effectiveness under low-resource settings.
@@ 4.3 Effects of Two-Stage Fine-tuning
We assess the impact of each stage in the two-stage LMM-RGCL fine-tuning process. As shown in Table 3(a) and Table 3(b), omitting either stage results in performance losses under both supervised and cross-dataset settings. Excluding Stage 1 results in the largest performance drop, particularly because the Large Multimodal Model (LMM) backbone remains frozen during Stage 2.
@@
LMM-RGCL 91.1 82.1 93.2 88.1
     w/o Stage 1 84.4 74.2 90.1 85.6
     w/o Stage 2 90.2 81.4 92.0 86.2
     Combined stages 88.9 77.8 90.2 83.4
@@
LMM-RGCL 77.1 69.3 88.8 81.7
     w/o Stage 1 72.0 62.1 84.9 78.1
     w/o Stage 2 74.4 66.7 86.3 78.7
     Combined stages 72.2 65.3 87.5 80.2
@@
Table 3: Ablation study of LMM-RGCL two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of Stage 1 Joint Multimodal Fine-tuning and Stage 2 RGCL Fine-tuning. For ‘combined stages’, we jointly optimize the three loss objectives from both stages in a single training process.
When stage 2 is omitted, the performance loss in supervised settings is less severe than in cross-dataset settings. We attribute this to the use of contrastive loss in Stage 2, which explicitly optimizes retrieval capabilities by aligning representations of semantically similar meme pairs. This alignment enhances robustness to distribution shifts in unseen datasets.
For ‘combined stages’, we jointly optimize the objectives for both stages:
$\mathcal{L}_{i}^{\textrm{Combined}}=\mathcal{L}_{i}^{RGCLL}+\mathcal{L}_{i}^{LR}+\mathcal{L}_{i}^{LM}.$ (9)
It yields suboptimal results, with performance drops of 2.2% AUC on HatefulMemes and 3.0% AUC on HarMeme under supervised settings. These results show that the two-stage fine-tuning approach effectively mitigates the optimization conflict between task adaptation in stage 1 and representation alignment in stage 2. Furthermore, since the LMM remains trainable under this combined loss objective, updating the encoded meme database incurs significantly higher computational costs compared to the two-stage fine-tuning approach, where the LMM is frozen in stage 2. This separation of fine-tuning stages enables more efficient training while preserving strong performance. Additional details on training time are provided in Appendix B.
For each stage, we perform an ablation study by removing individual loss objectives during fine-tuning. Our results indicate that each term is essential for optimal performance. Detailed results are provided in Appendix E.
@@ 4.4 Numbers of Shots and Neighbors
Figure 2 shows that as the number of nearest neighbors $K$ for RKC increases, the performance continues to increase for both AUC and accuracy, plateauing at around $K=20$. The consistent improvement in performance indicates that RKC trained with LMM-RGCL utilizes demonstration examples more effectively than the standard in-context learning framework.
Figure 2: Effects of increasing number of shots for in-context learning with pre-trained LMM and effects of increasing top K nearest neighbors for RKC trained with LMM-RGCL
@@ 4.5 Comparing Different Inference Modes
Table 4 compares Qwen2VL-7B fine-tuned with LMM-RGCL using the three classifiers. Our results indicate that under supervised settings, the differences among the three inference modes are minimal. However, under cross-dataset settings, there is a significant disparity in generalization performance. Notably, RKC outperforms both LMH and LRC, underscoring its superior effectiveness in handling out-of-domain examples.
HatefulMemes HarMeme
Inference Mode AUC Acc. AUC Acc.
LMH 90.2 81.9 92.8 88.0
LRC 91.1 82.1 93.2 88.1
RKC 90.8 81.8 93.2 88.0
(a) Supervised settings, see Table 1 for detailed settings
HatefulMemes HarMeme
Inference Mode AUC Acc. AUC Acc.
LMH 70.2 64.3 64.5 60.3
LRC 59.5 55.4 57.9 52.2
RKC 77.1 69.3 88.8 81.7
(b) Cross-dataset settings, see Table 2 for detailed settings
Table 4: Comparing different inference modes using LMM-RGCL fine-tuned Qwen2VL-7B. RKC shows much better out-of-domain generalization compared to other inference modes.
@@ 5 Conclusion
We propose LMM-RGCL, a two-stage fine-tuning framework designed to improve LMM performance on hateful meme classification, addressing the ineffectiveness of standard SFT. Our approach effectively improves both in-domain accuracy and cross-domain generalization. State-of-the-art performance across six meme classification datasets demonstrates the effectiveness of LMM-RGCL.
@@
Hateful meme detection systems, like LMM-RGCL, can be used to automatically detect hateful content online, contributing significantly to reducing online hate speech. By reducing hate speech, fostering safer digital environments, and supporting human content moderators, these systems can make a significant impact on online communication and safety. We believe these benefits are both substantial and essential in the broader effort to create a more secure and respectful digital space.
@@
Training large-scale models is computationally intensive and contribute to global warming due to heavy GPU/TPU usage. However, our approach mitigates this issue by fine-tuning LMMs using quantized LoRA, a parameter-efficient method. As a result, our system can be trained in under four hours on a single GPU, significantly reducing both training time and computational cost compared to full-scale LMM fine-tuning. Furthermore, since our method generalizes across different domains without requiring retraining, it further minimizes computational overhead.
@@ Appendix B Experiment Setup and Implementation Details
We conducted our experiments on a workstation equipped with NVIDIA RTX 3090 was used for the experiments. The full parameter fine-tuning experiments were carried out on 4 A100-80GB-SXM GPUs.
@@
The run time for LMM-RGCL two-stage fine-tuning on the HatefulMemes dataset is approximately 4 hours on a single RTX 3090, which is comparable to SFT.
@@
For full-parameter fine-tuning, training takes 6 hours on 4 A100-80G.
@@ B.1 LLaVA and Qwen2VL experiments
We freeze the vision module throughout fine-tuning, following the standard LMM fine-tuning protocol. For prompt formatting, we adhere to InstructBLIP Dai et al. (2023). For LLaVA few-shot experiments, since LLaVA is not explicitly trained to support in-context learning, we follow the procedure outlined by Zong et al. (2024) to enable few-shot learning on LLaVA. For fine-tuning LLaVA Liu et al. (2023b, a), we follow the original hyperparameters setting [^3] https://github.com/haotian-liu/LLaVA for fine-tuning on downstream tasks for both the SFT and LMM-RGCL stage 1 fine-tuning.
For Qwen2VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and LMM-RGCL stage 1 fine-tuning. For few-shot learning with Qwen2VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.

### v2

@@ 4.1 Comparing RA-HMD to Baseline Systems under Supervised Settings
13 RA-HMD + RKC 77.1 69.3 88.8 81.7 81.4 75.6 64.5 66.4 63.8 55.6 69.3 69.3
Table 2: Comparing out-of-domain meme classification performance under low-resource settings. For systems fine-tuned under cross-dataset settings, models are fine-tuned on HarMeme and evaluated on the HatefulMemes dataset. For the remaining evaluation datasets, models are fine-tuned on the HatefulMemes dataset. For LMMs, we compare the SFT models using zero-shot and in-context learning with the RA-HMD fine-tuned models using RKC. Few-shot examples (4-shot) and RKC examples are drawn from the training split of each evaluation dataset. #2 and #3 are taken from the original paper. Best performance is highlighted in bold; second-best is underlined.
@@
As shown in Table 1, RGCL (#3) achieves the highest performance among CLIP-based classifiers, surpassing standard fine-tuned CLIP (#1) by approximately 10% across multiple datasets. On 5 out of 6 datasets, RGCL performs better than, or on par with, all three SFT LMMs (#7, #11, #15).
@@
We compare the zero-shot (#5, #9, #13) and few-shot (#6, #10, #14) performance of the pre-trained LMMs. Our findings indicate that, in-context learning does not benefit meme classification, which is consistent with previous results Hee et al. (2024a); Huang et al. (2024). HarMeme is the only dataset where few-shot systems consistently outperform zero-shot systems. On Harm-P and MultiOFF, although the accuracies of zero-shot and few-shot remain comparable, the few-shot experiments yield a significant gain in F1 score. This improvement is due to a more balanced precision and recall after providing demonstration examples to the system.
Observation 3: RA-HMD outperforms all strong baseline systems across six datasets
Across six datasets and three LMMs, fine-tuning with RA-HMD significantly improves performance over SFT (Table 1: #7, #8; #11, #12; #15, #16). Statistical significance tests comparing RA-HMD and SFT further validate these results, with all p-values below 0.05 (see Appendix C). Notably, as indicated in #16, Qwen2VL-7B fine-tuned with RA-HMD outperforms VPD-PaLI-X-55B on HatefulMemes. Moreover, RA-HMD improves upon RGCL with gains of over 4% in AUC and 3% in accuracy on HatefulMemes. These gains show RA-HMD’s effectiveness in improving LMMs for meme classification over SFT.
4.2 Comparing RA-HMD with Baseline Systems under Low-Resource Settings
Online hate speech is constantly evolving, posing a challenge to systems as the distribution of memes encountered in the wild departs from that of the training data. To simulate real-world deployment constraints, we evaluate systems on out-of-domain examples under low-resource settings where gradient updates are prohibited and only demonstration examples are available Huang et al. (2024); Hee et al. (2024a); Cao et al. (2024).
@@ 4.2 Comparing RA-HMD with Baseline Systems under Low-Resource Settings
We compare RA-HMD fine-tuned LMM with the RKC against the following systems: SFT LMMs with zero-shot and few-shot prompting using LMH; GPT-4o OpenAI (2024a); specialized low-resource systems (LOREHM Huang et al. (2024), Mod-hate Cao et al. (2024)). For GPT-4o, we report results without prompt optimization for each dataset, as this setting assumes the hate type is not known in advance. Further discussion and comparison of GPT-4o results can be found in Appendix J.
@@
Observation 2: SFT LMMs with in-context learning is ineffective
As shown in Table 2 #6, #9 and #12, the few-shot approach remains similarly ineffective after LMMs are fine-tuned on different domains of hateful meme datasets, offering no significant gains over the SFT zero-shot models (Table 2 #5, #8, #11). In Section 4.5, we further analyze the effect of the number of shots in in-context learning and find that increasing the number of shots does not improve performance. Moreover, in Section 4.6, we show that in-context learning with RA-HMD consistently outperforms SFT models, demonstrating RA-HMD’s effectiveness in enhancing out-of-domain generalization.
Observation 3: RA-HMD fine-tuned LMMs with RKC inference mode outperforms baseline methods
RA-HMD fine-tuned LMMs using RKC outperform the baseline SFT LMMs in both zero-shot and few-shot settings under the same cross-dataset fine-tuning settings. Notably, RA-HMD trained Qwen2VL-7B with RKC improves over the baseline SFT few-shot model by 21.6% in AUC and 19.3% in accuracy on HarMeme (Table 2 #11-13). The ablation study in Section 4.5, which varies the number of top k for RKC, further demonstrates that RKC uses demonstration examples more effectively than the few-shot in-context learning framework. Moreover, as shown in Appendix F, applying RKC to both pretrained and SFT LMMs results in worse performance compared to RA-HMD, underscoring the effectiveness of our fine-tuning strategy.
Observation 4: RA-HMD trained LMMs with RKC inference outperform other low resource methods
When compared to other low-resource methods, our RA-HMD fine-tuned LLaVA-1.5-7B with RKC matches the performance of LOREHM on the HatefulMemes dataset. Notably, LOREHM uses a newer and larger LLaVA-1.6-34B within an agent-based framework. Furthermore, our method outperforms LOREHM by 8.2% in accuracy on HarMeme, highlighting our methods’ effectiveness under low-resource settings.
@@ 4.3 Effects of Two-Stage Fine-tuning
We assess the contribution of each stage within our two-stage RA-HMD fine-tuning process. As shown in Tables 3, omitting either stage leads to performance degradation in both supervised and cross-dataset settings, with Stage 1 contributing more substantial gains.
When only Stage 1 is applied and Stage 2 is omitted, the performance loss is less severe under supervised settings than in cross-dataset evaluations. We attribute this to the contrastive loss in Stage 2, which explicitly optimizes retrieval by aligning representations of semantically similar meme pairs, thereby enhancing robustness to distribution shifts in unseen datasets.
We also compared a variant where we jointly optimize the losses from both stages in a single training phase.
$\mathcal{L}_{i}^{\textrm{Combined}}=\mathcal{L}_{i}^{CL}+\mathcal{L}_{i}^{LR}+\mathcal{L}_{i}^{LM}.$ (9)
This combined training yields suboptimal results, demonstrating that the two-stage fine-tuning effectively resolves the optimization conflict between task adaptation (Stage 1) and representation alignment (Stage 2). Furthermore, since the LMM remains trainable throughout combined training, updating the encoded meme database incurs significantly higher computational costs compared to the two-stage fine-tuning approach, where the LMM is frozen in stage 2. This staged separation thus enables more efficient training while obtaining stronger performance.
In Appendix H, we also conduct ablation studies by removing individual loss terms within each stage and found that every loss component is essential.
@@
RA-HMD 91.1 82.1 93.2 88.1
w/ Stage 1 only 90.2 81.4 92.0 86.2
w/ Stage 2 only 84.4 74.2 90.1 85.6
w/ $\mathcal{L}_{i}^{\textrm{Combined}}$ 88.9 77.8 90.2 83.4
@@
RA-HMD 77.1 69.3 88.8 81.7
w/ Stage 1 only 74.4 66.7 86.3 78.7
w/ Stage 2 only 72.0 62.1 84.9 78.1
w/ $\mathcal{L}_{i}^{\textrm{Combined}}$ 72.2 65.3 87.5 80.2
@@
Table 3: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of Stage 1 and Stage 2 Fine-tuning. For $\mathcal{L}_{i}^{\textrm{Combined}}$, we jointly optimize the three loss objectives from both stages in a single training process as shown in Eq. 9.
4.4 Performance on General Vision-Language Benchmarks
Table 4 compares the pretrained Qwen2VL-2B with its SFT and RA-HMD variants, both fine-tuned on the HatefulMemes dataset, across three general vision-language benchmarks: MMMU Yue et al. (2023), SEED-Bench Li et al. (2023a), and GQA Hudson and Manning (2019). Evaluation settings and additional results are provided in Appendix I. The SFT model shows performance degradation across all three benchmarks, while RA-HMD maintains performance comparable to the pretrained model. These results indicate that RA-HMD robustly preserves the general vision-language capabilities of LMMs.
Model MMMU SEEDBench GQA
Qwen2VL-2B 40.2 72.7 60.4
+SFT 39.1 72.1 57.0
+RA-HMD 40.4 72.7 60.1
Table 4: Comparison of the pretrained, SFT, and RA-HMD Qwen2VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
@@ 4.5 Numbers of Shots and Neighbors
In Appendix E, we ablate the effects of varying the number of shots for few-shot in-context learning and varying the number of top K nearest neighbors for RKC. We find that merely adding more shots does not necessarily improve performance for in-context learning, aligning with the findings of Huang et al. (2024). In contrast, increasing $K$ in RKC leads to steady performance gains, with improvements plateauing around $K=20$. These results suggest that the RKC inference mode of RA-HMD makes more effective use of demonstration examples compared to in-context learning.
4.6 Comparing Out-of-Domain Generalization Across Model Variants
To further evaluate out-of-domain generalization across different model variants, we compare RA-HMD against both pretrained and SFT models in Appendix F, following the same protocol as in Table 2. Our results show that RA-HMD consistently outperforms both baselines under both the in-context learning framework and the retrieval-augmented KNN classifier (RKC), highlighting its robustness and effectiveness for generalizing to unseen meme distributions.
@@ 4.7 Comparing Different Inference Modes
We compare RA-HMD using three classifiers: LMH, LRC and RKC in Appendix G. Under supervised settings, the performance is similar. However, in cross-dataset scenarios, RKC outperforms both LMH and LRC, underscoring its superior effectiveness in handling out-of-domain examples.
4.8 Demonstration examples
In Appendix L, we present a case analysis comparing the classification results of RA-HMD and SFT. Appendix K provides example rationales generated by each model.
@@ 5 Conclusion
We propose RA-HMD, a robust adaptation framework for LMMs tailored for hateful meme classification. Our approach effectively improves both in-domain accuracy and out-of-domain generalization, achieving state-of-the-art results across six meme classification datasets while preserving the general vision-language capabilities of the underlying models.
@@
Hateful meme detection systems, like RA-HMD, can be used to automatically detect hateful content online, contributing significantly to reducing online hate speech. By reducing hate speech, fostering safer digital environments, and supporting human content moderators, these systems can make a significant impact on online communication and safety. We believe these benefits are both substantial and essential in the broader effort to create a more secure and respectful digital space.
@@
Training large-scale models is computationally intensive and contribute to global warming due to heavy GPU/TPU usage. However, our approach mitigates this issue by fine-tuning LMMs using quantized LoRA, a parameter-efficient method. As a result, our system can be trained in under four hours on a single consumer-grade GPU RTX3090, costing less than 1 USD, significantly reducing both training time and computational cost compared to full-scale LMM fine-tuning. Furthermore, since our method generalizes across different domains without requiring retraining, it further minimizes computational overhead.
@@ Appendix A Dataset details and statistics
For HatefulMemes, HarMeme, and MAMI, we report the Area Under the Receiver Operating Characteristic Curve (AUC) and Accuracy (Acc) in line with previous studies Kumar and Nandakumar (2022); Cao et al. (2023); Mei et al. (2024); Cao et al. (2024). For Harm-P, MultiOFF, and PrideMM, we report Accuracy and F1 score consistent with the literature Pramanick et al. (2021b); Mei et al. (2024); Shah et al. (2024); Lin et al. (2024).
@@ Appendix B Experiment Setup and Implementation Details
We conducted our experiments on a workstation equipped with NVIDIA RTX 3090 was used for the experiments.
@@
The run time for RA-HMD two-stage fine-tuning on the HatefulMemes dataset is approximately 4 hours on a single NVIDIA RTX 3090 GPU, and costs around 1 USD.
@@ B.1 LLaVA and Qwen2VL experiments
We freeze the vision module throughout fine-tuning, following the standard LMM fine-tuning protocol. For prompt formatting, we adhere to InstructBLIP Dai et al. (2023). For LLaVA few-shot experiments, since LLaVA is not explicitly trained to support in-context learning, we follow the procedure outlined by Zong et al. (2024) to enable few-shot learning on LLaVA. For fine-tuning LLaVA Liu et al. (2023b, a), we follow the original hyperparameters setting [^3] https://github.com/haotian-liu/LLaVA for fine-tuning on downstream tasks for both the SFT and RA-HMD stage 1 fine-tuning.
For Qwen2VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and RA-HMD stage 1 fine-tuning. For few-shot learning with Qwen2VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.